40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4818     Entry time: Tue Jun 14 18:12:34 2011
Author: Jamie, Kiwamu 
Type: Update 
Category: LSC 
Subject: LSC seems to be fully recovered 

We are now locking the arms reliably, with reasonable transmitted power.  We zeroed the LSC offsets with script, since they were apparently not being reset with either the overall burt restore or the arm restore scripts.

We have lost a bit of power through the mode cleaner.  However, we have opted not to tweak it up just yet, so that we don't have to realign to the arms.

ELOG V3.1.3-